# Enhancing Conversational Agents from Open-Source Large Language Models with Illocutionary Force and Document-Based Knowledge Retrieval

### v1

@@ 7 Conclusion and Future work
The results from this research unlock several potentials for the inclusion of communicative intent (illocutionary forces) which have been overlooked by computational linguistics in improving the responses given by conversational agents in an HCI setting. The illocutionary forces explored in this research originate from a dialogue domain making it very relevant as an additional linguistic feature layer which is not currently embedded in the context of transformer models. To improve on this work, there is a need to explore larger encoder-decoder models with larger parameters and explore the impact of repeating these experiments on a more robust GPU.

### v2

@@ 6 Conclusion and Future work
The results from this research unlock several potentials for the inclusion of communicative intent (speech acts) which have been overlooked by computational linguistics in improving the responses given by conversational agents in an HCI setting. The speech actss explored in this research originate from a dialogue domain making it very relevant as an additional linguistic feature layer which is not currently embedded in the context of transformer models. To improve on this work, there is a need to explore larger encoder-decoder models with larger parameters and explore the impact of repeating these experiments on a more robust GPU. The open-source can be found here  [^4] ConvoStem.